3 cards are chosen at random from a standard 52-card deck. What is the probability that they form a pair? (A 3-card hand is a pair if two of the cards match in rand but the third card is different.)
Answer by stanbon(75887)

3 cards are chosen at random from a standard 52-card deck. What is the probability that they form a pair? (A 3-card hand is a pair if two of the cards match in rand but the third card is different.)
---
Pick a type of card: 13 ways
Pick a pair of that type: 4C2 = 6 ways
----
Pick a 3rd card that is not that type: 48 ways
----
P(event as described) = [13*6*48]/52C3 = 3744/22100 = 0.1694
=================
Cheers,
Stan H.